Analytic Number Theory and Surrounding Areas

Date: 29 Sep. - 3 Oct. Place: RIMS (Research Institute of Mathematical Science) Kyoto, JAPAN Organizer: Shigeki AKIYAMA (Niigata Univ.)

Program

Monday 29 Sep.

10:00 – 10:50 Michio Ozeki (Yamagata Univ.)

Mass formula for the Jacobi weight enumerators of type II codes and some relationships of it with Jacobi forms

11:00 – 11:50 Winfried Kohnen (Univ. Heidelberg · Germany)

Estimating Fourier coefficients of Siegel modular forms

13:30 – 14:00 Takao Komatsu (Hirosaki Univ.)

From continued fractions to real numbers

14:00 – 14:30 Yukio Ohkubo (Inter. Univ. Kagoshima)
     On the discrepancy of the sequences (nα + f(n))
14:30 – 15:00 Rie Natsui (Keio Univ.)
      Principal convergents and mediant convergents

associated to α-continued fractions

15:30 – 16:00 Yumiko Ichihara (Waseda Univ. · JSPS)

The convergence of L-functions associated with Hilbert cusp form

16:00 – 16:30 Takaaki Tanaka (Keio Univ.)

Algebraic independence of certain power series associated with d-adic expansion of real numbers Tuesday 30 Sep.

9:00 – 9:50 Masanobu Kaneko (Kyushu Univ.)
     On modular forms satisfying a certain differential equation
10:00 – 10:50 Edward Burger (Williams College · USA)
      On a question of Davenport and a quantitative refinement

of the Markoff Spectrum

11:00 – 11:40 Shigeru Kanemitsu (Kinki Univ.), Yoshio Tanigawa (Nagoya Univ. ),

Haruo Tsukada (Kinki Univ.), Masami Yoshimoto (Nagoya Univ. · JSPS) An application of zeta functions

13:30 – 14:20 Noriko Hirata-Kohno (Nihon Univ.)
      Linear forms in p-adic logarithms and abc conjecture
14:30 – 15:00 Hidehiko Mishou (Nagoya Univ.), Hirofumi Nagoshi (Niigata Univ. · JSPS)

On value distribution of quadratic L-functions

15:30 – 16:00 Yuichi Kamiya (Nagoya Univ. · JSPS)
      On spectrums of certain harmonic functions

attached to the Riemann zeta-function

16:00 – 16:30 Takeshi Kurosawa (NTT SI Lab.)
      Arithmetic properties of reciprocal sums of binary recurrences
16:45 – 17:15 Chiharu Kaminishi (Keio Univ.)

Estimates of automorphic L-functions in the d-aspect

17:15 – 18:00 Jun’ichi Tamura
      A new approach to higher dimensional continued fractions

Wednesday 1 Oct.

9:00 – 9:50 Okazaki Ryotaro (Doshisha Univ)
      Counting integer points by Baker theory and basic math
10:00 – 10:50 Attila Pethő (Debrecen Univ. · Hungary)
      Generalized radix representations and dynamical systems
11:00 – 11:50 Yoichi Motohashi (Nihon Univ. )
      A Vista of Mean Zeta Values, Part 2
Afternoon  Free Time

Thursday 2 Oct.

9:00 – 9:50 Stéphane Louboutin (Univ. Caen · France)
      Effective bounds for Zeta and L-functions,

with applications to the Brauer-Siegel theorem

10:00 – 10:50 Gilles Lachaud (CNRS · France)
      Eisenstein series and the Riemann Hypothesis
11:00 – 11:30 Shin’ichiro Okada (Keio Univ.), Iekata Shiokawa (Keio Univ.)

< q, r > number system and algebraic independence

13:30 – 14:20 Shunji Ito (Kanazawa Univ.)

Purely periodic beta expansions with Pisot basis

14:30 – 15:00 Leo Murata (Meijigakuin Univ.), Koji Chinen (Osaka Inst. Tech.)

On a distribution property of the residual order of a(modp), IV

15:30 – 16:00 Yoshiyuki Kitaoka (Meijo Univ.), Michihiro Nozaki (ChitaHigashi High School)

A conjecture on decimal expansions of rational numbers

16:00 – 16:30 Asako Nakamura (Keio Univ.)

Dirichlet’s prime number theorem for P GL over function fields

17:00 – 17:30 Ryuuta Hashimoto (Nagoya Univ.)

On the continued fraction of roots of trinomials

17:30 – 18:00 Kenji Nagasaka (Hosei Univ.)
      Sound-image processing and applied systems

Friday 3 Oct.

10:00 – 10:50 Kohji Matsumoto (Nagoya Univ.)
     Mean values of automorphic L-functions attached to Doi-Naganuma and Ikeda lifts
11:00 – 11:50 Laurent Habsieger (CNRS · France)

Combinatorics and moments of L-function

13:30 – 14:00 Koichi Kawada (Iwate Univ.)
      On sums of cubes of primes
14:00 – 14:30 Masami Yoshimoto (Nagoya Univ. · JSPS)
      On the square mean of L(1, χ) with respect to characters
14:30 – 15:00 Hitoshi Nakada (Keio Univ.)
      On non-archimedean metric diophantine approximations
15:30 – 16:00 Yohei Tachiya (Keio Univ.)
      Irrationality of a certain Lambert series
16:00 – 16:30 Hideaki Ishikawa (Niigata Univ.)

On a difference between values of L(1/2 + it, χ_{j}) and L(1/2 + it, χ_{k})